# Leveraging Spreading Activation for Improved Document Retrieval in Knowledge-Graph-Based RAG Systems

### v1

@@ Abstract
Despite initial successes and a variety of architectures, retrieval-augmented generation (RAG) systems still struggle to reliably retrieve and connect the multi-step evidence required for complicated reasoning tasks. Most of the standard RAG frameworks regard all retrieved information as equally reliable, overlooking the varying credibility and interconnected nature of large textual corpora. GraphRAG approaches offer potential improvement to RAG systems by integrating knowledge graphs, which structure information into nodes and edges, capture entity relationships, and enable multi-step logical traversal. However, GraphRAG is not always an ideal solution as it depends on high-quality graph representations of the corpus, which requires either pre-existing knowledge graphs that are expensive to build and update, or automated graph construction pipelines that are often unreliable. Moreover, systems following this paradigm typically use large language models to guide graph traversal and evidence retrieval, leading to challenges similar to those encountered with standard RAG. In this paper, we propose a novel RAG framework that employs the spreading activation algorithm to retrieve information from a corpus of documents interconnected by automatically constructed knowledge graphs, thereby enhancing the performance of large language models on complex tasks such as multi-hop question answering. Experiments show that our method achieves better or comparable performance to iterative RAG methodologies, while also being easily integrable as a plug-and-play module with a wide range of RAG-based approaches. Combining our method with chain-of-thought iterative retrieval yields up to a 39% absolute gain in answer correctness compared to naive RAG, achieving these results with small open-weight language models and highlighting its effectiveness in resource-constrained settings.
@@ 1 Introduction
Initially proposed in [guu2020retrieval, lewis2020retrieval], Retrieval-Augmented Generation (RAG) has become a popular technique for improving the capabilities of Large Language Models (LLMs) across various tasks, including question answering and code generation. The basic idea of RAG is to provide LLM access to an external, easily updatable knowledge source by integrating it with the Information Retrieval (IR) component. This technique aims to improve the quality of LLM responses by reducing hallucinations and enhancing text creation accuracy and coherence. Numerous enhancements to the foundational RAG paradigm have been further proposed, ranging from innovative pipeline designs to improvements in individual components. Additionally, several surveys have been conducted to classify, decompose, and identify key elements within the myriad of proposed methodologies, helping track their evolution and highlighting promising future directions [zhao2024retrieval, fan2024survey, gao2023retrieval, gupta2024comprehensive].
@@
A particularly active research area explores the combination of RAG with graph-based systems that represent diversified and relational information. The goal of these integrations is to improve performance on complex tasks involving large, structured, and interconnected knowledge corpora, resulting in methodologies sometimes referred to as GraphRAG. Comprehensive surveys covering the background, components, downstream tasks, and industrial use cases of GraphRAG, as well as the technologies and evaluation methods used, are available in [peng2024graph, han2024retrieval].
@@
As mentioned in [peng2024graph], knowledge graph question answering (KGQA) is a key natural language processing task that has motivated the development of KG-based systems aimed at responding to user queries through structured reasoning over predefined knowledge graphs [zhang2021neural, lavrinovics2025knowledge, pan2024unifying]. The multi-hop question answering (MHQA) task is a related but broader concept, commonly applied in areas such as academic research, customer support, and financial or legal inquiries, where a comprehensive analysis of multiple textual documents is required. In this scenario, given a document corpus, the QA system generates responses to user queries by reasoning across multiple sources, involving sequential reasoning steps where responses at one step may rely on answers from prior steps. Several standard and graph-based RAG methodologies have been proposed to enhance the performance of LLMs on this specific task [zhang2023end, jeong2024adaptive, wang2024knowledge, jimenez2024hipporag]. Despite significant advancements, most of these approaches still rely primarily on standard or iterative retrieve-read-answer workflows, which are enhanced by optimized pipeline components, LLM decision-making, or LLM-guided graph traversal [singh2025agentic, fan2024survey]. However, hallucination and weak faithfulness remain unresolved problems in the context of MHQA. LLM-guided iterative retrieval may fail due to myopic knowledge exploration, retrieving context based on partial reasoning from previous steps, which can be incorrect or contain fabricated information. Similarly, one-step RAG systems with optimized retrieval components may fetch more useful evidence in the retrieved context, but they often introduce a lot of noise or fail to capture information from "bridge" documents whose entities or phrases are not mentioned in the input query. Thus, although adding improvements to vanilla RAG, the paradigm behind these advanced methods stays fundamentally the same, thereby overlooking powerful and well-established IR algorithms that can automatically identify relationships between documents and adapt retrieval for multi-hop query scenarios.
@@
Motivated by this gap, this paper proposes a novel GraphRAG framework that integrates Spreading Activation (SA), a method for searching associative networks originating in cognitive psychology. Beyond information retrieval [crestani1997application], SA has been successfully applied in several domains, including natural language processing [pollack1982natural, tsatsaronis2007word], robotics, and reinforcement learning [kono2019activation]. Specifically, we propose a RAG system that operates on a hybrid structure combining a knowledge graph with textual information from knowledge documents. It retrieves relevant documents from textual corpora by executing an SA algorithm through the knowledge graph, identifying and fetching chunks of text with the most pertinent information required to answer the multi-hop query, thereby capturing the reasoning structure and key relationships between entities mentioned in the query.
@@
To the best of our knowledge, only one prior paper has integrated RAG and SA-based methodologies [wu2025kg]. However, our approach differs in several key respects. First, the mentioned system relies on human-crafted knowledge graphs, whereas our pipeline includes an automated knowledge-graph construction phase. Second, it uses a prompted LLM to perform the SA procedure and expand the initially retrieved subgraph; by contrast, we perform SA automatically by exploring the knowledge graph in a breadth-first manner and spreading the activation based on edge weights assigned by an embedding model. Finally, the prior method fine-tunes the LLM to improve its instruction-following abilities, while our method requires no retraining. Overall, we believe our methodology provides an innovative direction for enhancing retrieval-augmented generation by leveraging a cognitively inspired mechanism that emphasizes associative relevance rather than surface-level similarity.
@@ 2.1 RAG and Multi-Hop Question Answering
Although traditional RAG systems enhance the capabilities of LLMs for both knowledge-intensive and standard NLP tasks, they encounter several critical issues that limit their performance when handling complex tasks. This is particularly evident in MHQA [mavi2024multi], where the system is required to reason across multiple pieces of information or documents to answer complicated questions. Traditional systems rely on simple, single-shot retrieval, which often fails to capture all necessary information or may even overload the LLM’s context window with unnecessary or incorrect data. Numerous additions are proposed through the evolution of RAG paradigms with the aim of overcoming naive "retrieve-then-generate" methodology, which, according to the survey [fan2024survey], can be classified into two categories: training-free and training-based approaches. Training-free approaches aim to optimize RAG pipelines through techniques such as prompt engineering, in-context learning, improved organization of data in the knowledge corpus, or agentic design [singh2025agentic]. In contrast, training-based methods rely on fine-tuning the retriever and generator components to boost the system’s performance on downstream tasks.
@@
Articles [wang2023query2doc, gao2023precise] propose methods that first generate plausible answer-like passages for the input query and then embed them using sparse or dense encoders to match semantically similar real documents from the knowledge base.
Papers [ma2023query, fu2021decomposing] suggest that complex, implicit, or multi-hop queries should be rewritten, disambiguated, or decomposed before retrieval to optimize the fetched context.
@@
In [trivedi2022interleaving] authors propose a RAG system integrated with Chain‑of‑thought (CoT), a prompting technique, where each CoT step becomes the query for the next retrieval. Similarly,  [kim2023tree] proposes system based on Tree‑of‑Thoughts (ToT) structured prompting framework; starting from an underspecified root question, the model generates candidate clarification subquestions, prunes uninformative branches, and for each surviving node runs a retrieval round targeted to that clarification, so that by the time the leaves are answered the system has accumulated disjoint, high-precision evidence for answering the original query.
Few papers propose integrating advanced RAG mechanisms with reflection modules that allow LLMs to determine the most suitable retrieval strategy based on the input query [jeong2024adaptive, wang2023self, asai2024self, crag]. By reflecting on the question, the LLM can choose whether to perform simple retrieval, engage in iterative retrieval steps, or rely on its own internal knowledge. Some systems can dynamically guide the direction and number of retrieval steps, evaluate the relevance of retrieved content [asai2024self], or make use of external tools such as online search [crag].
@@ 2.2 GraphRAG
The GraphRAG paradigm [peng2024graph, han2024retrieval] aims to extend RAG systems by leveraging the structured relationships and hierarchical organization of information in graph data to improve multi-hop reasoning and contextual understanding.
One of the first successful GraphRAG systems, introduced in [edge2024local], was designed for Query-Focused Summarization (QFS), where, provided a query and a document corpus, the system seeks to compress the contents of multiple documents, gathering the identified themes and providing a unified summary of the entire document corpus. The proposed approach focuses on the connection between the summarization task and the inherent modularity of graphs, which allows partitioning of graphs into communities. LLM is first instructed to create a knowledge graph from source texts by extracting entities and relationships between them. The system then uses the Leiden community detection algorithm to partition this graph into hierarchical communities. Finally, each community is summarized separately, and the summaries are combined to produce a final, comprehensive response to the user’s question.
@@
LightRAG proposed in [guo2024lightrag] firstly builds a lightweight knowledge graph over a document corpus by using an LLM to extract entities and relations from chunks. For each node and edge, it then constructs a key–value entry, where the key is a short textual label and the value is a concise summary derived from the source documents. At query time, another language model analyses the input and produces two sets of query terms, one focusing on concrete objects and relations, and another capturing broader topics. Terms from the first set are embedded and used for vector search over keys tied to entity nodes, while the broader theme terms are used to match keys associated with relation edges. The retrieved graph elements are expanded by one hop to include their neighbors, then the corresponding value texts are combined into a compact context that is passed to a generative model to produce the final answer.
In [g-retriever], the authors present G-Retriever, a framework that leverages the integration of graph neural networks and LLMs. The system efficiently responds to user queries by retrieving relevant subgraphs using a Prize-Collecting Steiner Tree algorithm and translating graph data to a textual representation. The graph encoder and projection layer, which are in charge of aligning subgraph information with the language model’s hidden space, are fine-tuned via backpropagation based on the difference between the generated answer and the ground truth. This technique is evaluated on different tasks, ranging from commonsense reasoning and scene comprehension to knowledge-based question answering.
Another design for an MHQA pipeline is described in [jimenez2024hipporag], which includes creating a schemaless knowledge graph with instruction-tuned LLM. When a query is received, the language model extracts the key named entities from the query, which are then mapped to their associated nodes in the knowledge graph using similarity scores calculated by the retrieval encoder. These nodes serve as seeds for a graph search using the Personalized PageRank algorithm, which propagates activation through the network to identify nodes that may be associated with the provided question.
@@ 2.3 Spreading Activation
Paper [quillian1967word] proposed that human semantic memory is organized as a network and was one of the first to implement a semantic network model of knowledge in a computer. In this framework, concepts and their properties were modeled as nodes in a network connected by links encoding semantic relationships, arranged hierarchically so that more specific concepts point to more general ones. Activation spreading was introduced as a mechanism to search this network in order to verify factual sentences (e.g., canary is a bird). This was achieved by activating a concept and letting that activation move through the network until it intersected with activation from another concept. Later in [collins1969retrieval] the computational model was tested against human performance in sentence–verification tasks, where people’s response times to statements at different hierarchical levels (e.g. “a canary is a bird” vs. “a canary is an animal”) were compared to the model’s predictions, showing that longer network traversals correlate with slower verification times. This model was then modified in [collins1975spreading, anderson1983spreading] to account for varying association strengths and decay of activation over distance, dropping the requirement that the network is strictly hierarchical. The spreading mechanism was generalized so that activation spreads from a currently active concept to all linked concepts, in parallel, with the strength that depends on link weight and fades with distance.
@@
In [salton1988use], the authors argued that SA can be implemented in IR systems to overcome the limitations of standard TF-IDF-based approaches, particularly when the user’s wording does not match the wording of the documents and when relevant documents are only indirectly connected to the query. They proposed a system using network representation of the document corpus, where both documents and index terms are modeled as nodes, interconnected by various types of links: term–document links modeling the presence of a term in a document, term–term links connecting terms that frequently co-occur or are semantically related according to a thesaurus, and document–document links connecting documents that share co-occurring terms. In this setting, the SA procedure starts by taking the query terms or documents retrieved by the normal vector-space search and assigning them an initial activation value. The activation then propagates from every active node to its neighbors, with the amount passed along each link scaled by the link’s strength, and nodes whose accumulated activation exceeds a specified threshold are finally retrieved. However, the authors implemented and evaluated only a simplified version of the described procedure due to the computational costs of a full implementation at that time. They concluded that the spreading activation procedure is a useful complementary module to standard IR systems, but it needs to be constrained by limiting the number of steps to just a few hops from the original query nodes and by using a decay factor smaller than one. Later works expanded criticism of unconstrained SA along the same lines; In [berthold2009pure], the authors showed that the pure SA algorithm, where activation propagates over all outgoing links with no limits on hops, fan-out, or thresholds, converges to a fixed state and retrieval results become nearly independent of the input query. This underscores that heuristic constraints in applied SA-based systems are not just pragmatic features but essential for the retrieval of query‐dependent results.
@@
More recent applications in IR and recommendation systems, therefore, employ adapted, task-driven variants of SA, where the spread is guided by the user’s information need or context and then combined with conventional ranking steps. In [ngo2014discovering], the authors propose a semantic text search method based on SA, which, given an input query, performs named entity recognition and relationship extraction, then maps the extracted elements to an ontology to represent the query as a concise structured pattern. The method then spreads activation only to concepts that match the relationships actually expressed in the query, as well as to closely related information such as alternative names or parent categories. Activated concepts are retrieved and appended to the original query to form an expanded query, which is then executed by a classical sparse-vector search engine, allowing ordinary document ranking to exploit both the user’s words and the newly discovered latent concepts.
@@
Article [blanco2010adapting] describes a semantic, knowledge-based recommender built around an ontology that stores domain concepts, user interests, and item descriptors, and then runs a controlled spreading of activation from the user profile toward semantically adjacent items so that recommendations reflect both explicit preferences and latent relations present in the knowledge base. The system routes the activation only through specific, typed links in the ontology, assigning weight to each link based on its type and finally ranking associated items according to their activation values and standard recommendation filters. Similarly, [papneja2018context] extends ontology-based SA with contextual cues, such as current task or situational attributes, so that activation is fired not only from the long-term user profile but also from context nodes. Activation propagates only along meaningful ontology links, with decay and link weights, thereby avoiding uncontrolled diffusion through the network.

### v2

@@ Abstract
Despite initial successes and a variety of architectures, retrieval-augmented generation (RAG) systems still struggle to reliably retrieve and connect the multi-step evidence required for complicated reasoning tasks. Most of the standard RAG frameworks regard all retrieved information as equally reliable, overlooking the varying credibility and interconnected nature of large textual corpora. GraphRAG approaches offer potential improvement to RAG systems by integrating knowledge graphs, which structure information into nodes and edges, capture entity relationships, and enable multi-step logical traversal. However, GraphRAG is not always an ideal solution, as it depends on high-quality graph representations of the corpus. Such representations usually rely on manually curated knowledge graphs, which are costly to construct and update, or on automated graph-construction pipelines that are often unreliable. Moreover, systems following this paradigm typically use large language models (LLMs) to guide graph traversal and evidence retrieval, leading to challenges similar to those encountered with standard RAG. In this paper, we propose a novel RAG framework that uses a spreading activation algorithm to retrieve information from a corpus of documents connected by an automatically constructed heterogeneous knowledge graph. This approach reduces reliance on semantic knowledge graphs, which are often incomplete due to information loss during information extraction, avoids LLM-guided graph traversal, and improves performance on multi-hop question answering. Experiments show that our method achieves better or comparable performance than several state-of-the-art RAG methods and can be integrated as a plug-and-play module with different iterative RAG pipelines. When combined with chain-of-thought iterative retrieval, it yields up to a 39% absolute improvement in answer correctness over naive RAG, while achieving these results with small open-weight language models, supporting its use in settings with limited computational resources. Code used to run the experiments is available at https://github.com/jomibg/sa-rag
@@ 1 Introduction
Initially proposed in [17, 26], Retrieval-Augmented Generation (RAG) has become a popular technique for improving the capabilities of Large Language Models (LLMs) across various tasks, including question answering and code generation. The basic idea of RAG is to provide LLM access to an external, easily updatable knowledge source by integrating it with the Information Retrieval (IR) component. This technique aims to improve the quality of LLM responses by reducing hallucinations and enhancing text creation accuracy and coherence. Numerous enhancements to the foundational RAG paradigm have been further proposed, ranging from innovative pipeline designs to improvements in individual components. Additionally, several surveys have been conducted to classify, decompose, and identify key elements within the myriad of proposed methodologies, helping track their evolution and highlighting promising future directions [50, 10, 13, 15].
@@
A particularly active research area explores the combination of RAG with graph-based systems that represent diversified and relational information. The goal of these integrations is to improve performance on complex tasks involving large, structured, and interconnected knowledge corpora, resulting in methodologies sometimes referred to as GraphRAG. Comprehensive surveys covering the background, components, downstream tasks, and industrial use cases of GraphRAG, as well as the technologies and evaluation methods used, are available in [33, 18].
@@
As mentioned in [33], knowledge graph question answering (KGQA) is a key natural language processing task that has motivated the development of KG-based systems aimed at responding to user queries through structured reasoning over predefined knowledge graphs [49, 25, 31]. The multi-hop question answering (MHQA) task is a related but broader concept, commonly applied in areas such as academic research, customer support, and financial or legal inquiries, where a comprehensive analysis of multiple textual documents is required. In this scenario, given a document corpus, the QA system generates responses to user queries by reasoning across multiple sources, involving sequential reasoning steps where responses at one step may rely on answers from prior steps. Several standard and graph-based RAG methodologies have been proposed to enhance the performance of LLMs on this specific task [48, 21, 44, 22]. Despite significant advancements, most of these approaches still rely primarily on standard or iterative retrieve-read-answer workflows, which are enhanced by optimized pipeline components, LLM decision-making, or LLM-guided graph traversal [37, 10]. However, hallucination and weak faithfulness remain unresolved problems in the context of MHQA. LLM-guided iterative retrieval may fail due to myopic knowledge exploration, retrieving context based on partial reasoning from previous steps, which can be incorrect or contain fabricated information. Similarly, one-step RAG systems with optimized retrieval components may fetch more useful evidence in the retrieved context, but they often introduce a lot of noise or fail to capture information from "bridge" documents whose entities or phrases are not mentioned in the input query. Thus, although adding improvements to vanilla RAG, the paradigm behind these advanced methods stays fundamentally the same, thereby overlooking powerful and well-established IR algorithms that can automatically identify relationships between documents and adapt retrieval for multi-hop query scenarios.
@@
Motivated by this gap, this paper proposes a novel GraphRAG framework that integrates Spreading Activation (SA), a search method for associative networks originally developed in cognitive psychology. Beyond information retrieval [8], SA has been successfully applied in several domains, including natural language processing [34, 41], robotics, and reinforcement learning [24]. Specifically, we propose a RAG system that operates on a hybrid structure combining a knowledge graph with textual information from knowledge documents. It retrieves relevant documents from textual corpora by executing an SA algorithm through the knowledge graph, identifying and fetching chunks of text with the most pertinent information required to answer the multi-hop query, thereby capturing the reasoning structure and key relationships between entities mentioned in the query.
@@
To the best of our knowledge, only one prior paper has integrated RAG and SA-based methodologies [45]. However, our approach differs in several key respects. First, the mentioned system relies on human-crafted knowledge graphs, whereas our pipeline includes an automated knowledge-graph construction phase. Second, it uses a prompted LLM to perform the SA procedure and expand the initially retrieved subgraph; by contrast, we perform SA automatically by exploring the knowledge graph in a breadth-first manner and spreading the activation based on edge weights assigned by an embedding model. Finally, the prior method fine-tunes the LLM to improve its instruction-following abilities, while our method requires no retraining. Overall, we believe our methodology provides an innovative direction for enhancing retrieval-augmented generation by leveraging a cognitively inspired mechanism that emphasizes associative relevance rather than surface-level similarity.
@@ 2.1 RAG and Multi-Hop Question Answering
Although traditional RAG systems enhance the capabilities of LLMs for both knowledge-intensive and standard NLP tasks, they encounter several critical issues that limit their performance when handling complex tasks. This is particularly evident in MHQA [28], where the system is required to reason across multiple pieces of information or documents to answer complicated questions. Traditional systems rely on simple, single-shot retrieval, which often fails to capture all necessary information or may even overload the LLM’s context window with unnecessary or incorrect data. Numerous additions are proposed through the evolution of RAG paradigms with the aim of overcoming naive "retrieve-then-generate" methodology, which, according to the survey [10], can be classified into two categories: training-free and training-based approaches. Training-free approaches aim to optimize RAG pipelines through techniques such as prompt engineering, in-context learning, improved organization of data in the knowledge corpus, or agentic design [37]. In contrast, training-based methods rely on fine-tuning the retriever and generator components to boost the system’s performance on downstream tasks.
@@
Articles [42, 12] propose methods that first generate plausible answer-like passages for the input query and then embed them using sparse or dense encoders to match semantically similar real documents from the knowledge base.
Papers [27, 11] suggest that complex, implicit, or multi-hop queries should be rewritten, disambiguated, or decomposed before retrieval to optimize the fetched context.
@@
In [39] authors propose a RAG system integrated with Chain‑of‑thought (CoT), a prompting technique, where each CoT step becomes the query for the next retrieval. Similarly,  [23] proposes system based on Tree‑of‑Thoughts (ToT) structured prompting framework; starting from an underspecified root question, the model generates candidate clarification subquestions, prunes uninformative branches, and for each surviving node runs a retrieval round targeted to that clarification, so that by the time the leaves are answered the system has accumulated disjoint, high-precision evidence for answering the original query.
Few papers propose integrating advanced RAG mechanisms with reflection modules that allow LLMs to determine the most suitable retrieval strategy based on the input query [21, 43, 2, 46]. By reflecting on the question, the LLM can choose whether to perform simple retrieval, engage in iterative retrieval steps, or rely on its own internal knowledge. Some systems can dynamically guide the direction and number of retrieval steps, evaluate the relevance of retrieved content [2], or make use of external tools such as online search [46].
@@ 2.2 GraphRAG
The GraphRAG paradigm [33, 18] aims to extend RAG systems by leveraging the structured relationships and hierarchical organization of information in graph data to improve multi-hop reasoning and contextual understanding.
One of the first successful GraphRAG systems, introduced in [9], was designed for Query-Focused Summarization (QFS), where, provided a query and a document corpus, the system seeks to compress the contents of multiple documents, gathering the identified themes and providing a unified summary of the entire document corpus. The proposed approach focuses on the connection between the summarization task and the inherent modularity of graphs, which allows partitioning of graphs into communities. LLM is first instructed to create a knowledge graph from source texts by extracting entities and relationships between them. The system then uses the Leiden community detection algorithm to partition this graph into hierarchical communities. Finally, each community is summarized separately, and the summaries are combined to produce a final, comprehensive response to the user’s question.
@@
LightRAG proposed in [14] firstly builds a lightweight knowledge graph over a document corpus by using an LLM to extract entities and relations from chunks. For each node and edge, it then constructs a key–value entry, where the key is a short textual label and the value is a concise summary derived from the source documents. At query time, another language model analyses the input and produces two sets of query terms, one focusing on concrete objects and relations, and another capturing broader topics. Terms from the first set are embedded and used for vector search over keys tied to entity nodes, while the broader theme terms are used to match keys associated with relation edges. The retrieved graph elements are expanded by one hop to include their neighbors, then the corresponding value texts are combined into a compact context that is passed to a generative model to produce the final answer.
In [19], the authors present G-Retriever, a framework that leverages the integration of graph neural networks and LLMs. The system efficiently responds to user queries by retrieving relevant subgraphs using a Prize-Collecting Steiner Tree algorithm and translating graph data to a textual representation. The graph encoder and projection layer, which are in charge of aligning subgraph information with the language model’s hidden space, are fine-tuned via backpropagation based on the difference between the generated answer and the ground truth. This technique is evaluated on different tasks, ranging from commonsense reasoning and scene comprehension to knowledge-based question answering.
@@
Another MHQA pipeline is presented in [22, 16]. Their proposed approach, HippoRAG, creates a schemaless knowledge graph with instruction-tuned LLM. When a query is received, the language model extracts the key named entities from the query, which are then mapped to their associated nodes in the knowledge graph using similarity scores calculated by the retrieval encoder. These nodes serve as seeds for a graph search using the Personalized PageRank algorithm, which propagates activation through the network to identify nodes that may be associated with the provided question. Since this approach is similar to our solution in its use of a knowledge graph to retrieve relevant textual passages, we adopt it as a natural baseline in our experiments.
@@ 2.3 Spreading Activation
Paper [35] proposed that human semantic memory is organized as a network and was one of the first to implement a semantic network model of knowledge in a computer. In this framework, concepts and their properties were modeled as nodes in a network connected by links encoding semantic relationships, arranged hierarchically so that more specific concepts point to more general ones. Activation spreading was introduced as a mechanism to search this network in order to verify factual sentences (e.g., canary is a bird). This was achieved by activating a concept and letting that activation move through the network until it intersected with activation from another concept. Later in [7] the computational model was tested against human performance in sentence–verification tasks, where people’s response times to statements at different hierarchical levels (e.g. “a canary is a bird” vs. “a canary is an animal”) were compared to the model’s predictions, showing that longer network traversals correlate with slower verification times. This model was then modified in [6, 1] to account for varying association strengths and decay of activation over distance, dropping the requirement that the network is strictly hierarchical. The spreading mechanism was generalized so that activation spreads from a currently active concept to all linked concepts, in parallel, with the strength that depends on link weight and fades with distance.
@@
In [36], the authors argued that SA can be implemented in IR systems to overcome the limitations of standard TF-IDF-based approaches, particularly when the user’s wording does not match the wording of the documents and when relevant documents are only indirectly connected to the query. They proposed a system using network representation of the document corpus, where both documents and index terms are modeled as nodes, interconnected by various types of links: term–document links modeling the presence of a term in a document, term–term links connecting terms that frequently co-occur or are semantically related according to a thesaurus, and document–document links connecting documents that share co-occurring terms. In this setting, the SA procedure starts by taking the query terms or documents retrieved by the normal vector-space search and assigning them an initial activation value. The activation then propagates from every active node to its neighbors, with the amount passed along each link scaled by the link’s strength, and nodes whose accumulated activation exceeds a specified threshold are finally retrieved. However, the authors implemented and evaluated only a simplified version of the described procedure due to the computational costs of a full implementation at that time. They concluded that the spreading activation procedure is a useful complementary module to standard IR systems, but it needs to be constrained by limiting the number of steps to just a few hops from the original query nodes and by using a decay factor smaller than one. Later works expanded criticism of unconstrained SA along the same lines; In [3], the authors showed that the pure SA algorithm, where activation propagates over all outgoing links with no limits on hops, fan-out, or thresholds, converges to a fixed state and retrieval results become nearly independent of the input query. This underscores that heuristic constraints in applied SA-based systems are not just pragmatic features but essential for the retrieval of query‐dependent results.
@@
More recent applications in IR and recommendation systems, therefore, employ adapted, task-driven variants of SA, where the spread is guided by the user’s information need or context and then combined with conventional ranking steps. In [30], the authors propose a semantic text search method based on SA, which, given an input query, performs named entity recognition and relationship extraction, then maps the extracted elements to an ontology to represent the query as a concise structured pattern. The method then spreads activation only to concepts that match the relationships actually expressed in the query, as well as to closely related information such as alternative names or parent categories. Activated concepts are retrieved and appended to the original query to form an expanded query, which is then executed by a classical sparse-vector search engine, allowing ordinary document ranking to exploit both the user’s words and the newly discovered latent concepts.
@@
Article [4] describes a semantic, knowledge-based recommender built around an ontology that stores domain concepts, user interests, and item descriptors, and then runs a controlled spreading of activation from the user profile toward semantically adjacent items so that recommendations reflect both explicit preferences and latent relations present in the knowledge base. The system routes the activation only through specific, typed links in the ontology, assigning weight to each link based on its type and finally ranking associated items according to their activation values and standard recommendation filters. Similarly, [32] extends ontology-based SA with contextual cues, such as current task or situational attributes, so that activation is fired not only from the long-term user profile but also from context nodes. Activation propagates only along meaningful ontology links, with decay and link weights, thereby avoiding uncontrolled diffusion through the network.